OpenStudy (anonymous):

For a real number a and non-empty subset of reals B, define: a + B = { a + b : b is in B }. Show that if B is bounded above, then sup( a + B ) = a + sup B What I have so far: since B is bounded, ∀b∈B, b ≤ m, for some real number m ∀b∈B, a + b ≤ a + m. Since a + b is an arbitrary element in a + B, a + B is bounded above by a + m. Since R has the least upper bound properties, both a + B and B has a supremum. --------- The part I'm currently having trouble with is sup(a + B) =a + sup(B) My attempt so far is, ∀b∈B, b ≤ sup(B) by property of supremum. ∀b∈B, a + b ≤ a + sup(B) hence, sup(a+B) ≤ a + sup(B)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

For every \( \epsilon > 0\) ther is a \( b_\epsilon \) suich that \[ b_\epsilon > \sup(B) -\epsilon\\ \sup(a+B) \ge a + b_\epsilon > a +\sup(B) -\epsilon \] Since this is true for every \(\epsilon >0\), then we are done

OpenStudy (anonymous):

\(b_\epsilon \in B\)

OpenStudy (anonymous):

how is sup(a + B) >/ a + b_e ?

OpenStudy (anonymous):

What I understand so far is, sup(B) - e is not an upper bound of B. This means there is an element b_e in B such that sup(B) - e < b_e <= sup(B)

OpenStudy (anonymous):

adding a to the inequality gives a + sup(B) - e <a + b_e <= a + sup(B)

OpenStudy (anonymous):

oh wait, so a + b_e is some element of a + B, so by definition of sup, a + b_e <= sup(a + B). so we have: a + sup(B) - e < a + b_e <= sup(a + B)

OpenStudy (anonymous):

ok, so now the question is how does that imply that a + sup(B) <= sup(a + B) ??

OpenStudy (anonymous):

@AngusV Openstudy is not showing text symbols properly. And yeah, i hate proving the obvious too :/

OpenStudy (anonymous):

I hate questions like these. It's stupidly obvious that it is so. The question translates to: "If you take a number from a set of numbers and add it to another number, then you can't have a result greater than the number you've added + the largest number from the set." I'm trying to think of a way not to complicate things beyond that level. Your reasoning is perfectly correct because "a" is a constant real number outside B / the set of numbers. If a belongs to B and b belongs to B - then yeah, sup (a+b) =/= a+b. Don't overthink it, just write that ∀b∈B,b≤sup(B) because of sup(B) property of whatever - and then add "a" to both sides so that a+b≤a+sup(B) , where a∈R Since b is defined as any one element from B, it's like saying a+B≤a+sup(B) If your teacher has problems understanding why 2+3=3+2=5, I'm sorry for him/her and his/her degree.

OpenStudy (anonymous):

Gods fluttering damn this piece of pellet

OpenStudy (anonymous):

@AngusV I think you can just refresh the page, and all symbols will show up properly

OpenStudy (anonymous):

Oh !

OpenStudy (anonymous):

So yeah, in essence you can say it's due to "the property of inequations" or whatever.

OpenStudy (anonymous):

That if a < b then a+x < b+x

OpenStudy (anonymous):

thank you both @AngusV and @eliassaab . I appreciate your time :)